A Bibliography of Publications of Pieter W. Hemker

Pieter W. Hemker CWI

NL 1098 SJ Amsterdam The Netherlands

Tel: ?n/a? FAX: + 31 - 20 - 592 4199

E-mail: pieth@cwi.nl (Internet)

05 February 2018 Version 1.30

Abstract

This bibliography records publications of Pieter W. Hemker.

Title word cross-reference

2 [Hem82a, HM91]. 3 [Hem95a, Hem95b, Hem96a, HdZ97, Hem97, KH97]. D [KH92b]. ϵ [HSS97a, HSS99a, HSS00]. H [HHA09]. hp [HL08]. n [TH94].

-adaptive [HL08]. -component [TH94]. -D [Hem82a]. -Dimensional [HdZ97]. -transform [HHA09]. -Uniform [HSS97a, HSS99a, HSS00].

2D [MH90].

3-Dimensional [Hd93a, Hd93b]. 3D [Hem92, HKN97, HKN98, KHE97, NH97, NH98].

4 [Hem96a].

65 [HOS03].

8th [WHO08].

96c [Hem96a].

aan [Hem90a, Hem96b]. aanvaarding [Hem90a]. Acceleration [DHH^{+}91, HSS01b]. Accuracy [HSS97a, Hem87, HSS99b, HSS00, HSS02b, KSS04]. accurate [Hem82a, Hem82d, HSS01b, HSS01a, HSS02a, HSS02c, HSS03a, HSS03b, KH91c]. activity [HEH^{+}67, HH09, HH13]. actuator [ELE^{+}06]. acute [DHH^{+}02]. adapted [HP93]. Adaptive [DHH^{+}91, HS94b, HKL^{+}97b, Hem00, VH95, Hem80d, HvdME90, HM91, HS01, HL08, vdMHE90, NH00, SSH04, VKH06, vdMHKM93]. Adv. [Hem96a]. affinity [RTH^{+}11]. ALGOL [CHvS72, Hem71a, Hem73a, Hem77a, HW79, Hem81c]. algorithm [Hem80d]. Algorithms [Hem84c, DH90a, DH90b, Hem80b, Hem82f, Hem84b]. ambt [Hem90a]. Amsterdam [Hem90a]. Analysis [DH90a, DH90b, DH95, HM79, Hem80a, HHvR03b, HvR04, HHvR04, vRH05]. antibody [RTH^{+}11]. antibody-antigen [RTH^{+}11]. antigen [RTH^{+}11]. Application [HHH00, Hem00, NH00, vdMHKM93]. applied [ELE^{+}06, dH79]. approach [BHS84, Hem88a, HM91, HHvR03b, MH90]. approaches [HJ87, NH97, NH98]. Approximate [TH94, Hem82e]. Approximation [HS92, HS93, HS94a, HP96, HP97, TH91]. Approximations [FHS95, FHS96a, FHS96b, FHS96c]. Arising [HSS97b]. Aspects [HHL65, HH68, HSd80, Hem83b]. assay [HHv^{+}71]. assessment [RTH^{+}11]. autonomous [HWD84].

bases [HP93]. Basic [HHL65]. BASIS [HvdME90]. BASIS3 [Hd93a, Hd93b, HdZ97]. begin [CHvS72]. begin- [CHvS72]. behaviour [HD93c]. bias [Hem82a]. bibliography [Hem80c]. bij [Hem90a]. Bijzonder [Hem90a]. binding [RTH^{+}11]. bioassay [RTH^{+}11]. biological [HdH06]. Blood [HHL65, HH68, HdH06]. bouds [dH79]. Boundary [FHS95, FHS96a, Hem77b, HS94b, FHST95, FHS96b, FHS96c, dH79, Hem74a, HSd80, HSS02c, HHvR03a, HSS03a]. boundary-value [dH79, HSd80].

C1 [DHH^{+}02]. C1-inhibitor [DHH^{+}02]. calibration [HHA09]. cascades [HH69]. cases [HP93]. CFD [HK97, KH97, KHE97]. chemical [HDH05]. circle [Hem74b]. Class [HS94b, HST02, SSH04]. Clotting [HHL65, HH68, WHH06]. coagulation [HDH05, HdH06]. Coarsened [KHE97]. Coarsening [KHd97]. code [HWD84, vdMHE90]. codes [HSd80]. Colloquium [BHPv76, BDHv75, DHv72]. Comparing [HSd80]. comparison [Hem82f, RTH^{+}11]. competitive [HH68]. component [TH91, TH94]. composed [HST02]. Compressible [VH95, DHK93]. Comput [Hem96a]. Computation [Hem86a, HHv^{+}73, Hem96c, KH91b, KH92a]. computational [DHKL94]. Computations [KHd97, HvdME90, HSS01b, KdH88, KH91a, KH91c, KH92b, KvBH^{+}03]. concentrated [SSH97]. concentrations [HHA09]. Condition [HS92, HS93, HS94a]. conditions [FHST95, HSS02c, HHvR03a, HSS03a]. Conference [WHO08]. constants [RTH^{+}11]. Continuation [PWBH81]. Contributions [HW94a]. control [HDH05]. Convection [HSS97b, Hem82d, HSS99a, HSS01b, HSS01a, HSS02a, HSS02c, HSS02b, HST02, HSS03a, HSS03b, SSH04, VKH06, vRH05]. Convection-Diffusion [HSS97b, HSS99a, HSS01b, HSS02a, HSS02c, HST02, HSS03a, HSS03b, SSH04, VKH06, vRH05]. Convergence [DH95, HD93c, NH97, NH98, DH90a, DH90b]. converting [HEH^{+}67]. corrected [HH09, HH13]. Correction [Hem82b, BHS84, DH90a, DH90b, DH95, EH05, ELH08a, ELH08b, Hem82c, Hem82d, Hem82e, HD83, Hem83c, Hem84a, Hem86b, HK88a, HK88b, HK88c, HD93c, HK94, HK95a, HK95b, HS95, HSS97c, HSS98, HSS99a, HSS02a, HSS03b, KH91c]. Corrigendum [Hem96a]. course [HH09, HH13]. Criteria [Hem77a]. cubature [Hem73b]. curved [HL08].

D [Hem96a, Hem82a, HM91, Hem95a, Hem95b, Hem97, KH97]. D-problems [Hem96a, Hem95b, Hem97]. Damped [KH91a]. Data [Hd93a, Hd93b, HdZ97, HvdME90, SSH97]. decomposition [FHST95, Hem80b, KSS04]. Decoupled [DHH^{+}91]. deel [BDHv75, DHv72]. Defect [Hem82b, HD83, Hem86b, HK88a, HK88b, HK94, HK95a, HK95b, HSS02a, BHS84, DH90a, DH90b, DH95, EH05, ELH08a, ELH08b, Hem82c, Hem82d, Hem82e, Hem83c, Hem84a, HK88c, HD93c, HS95, HSS97c, HSS98, HSS99a, KH91c]. defect-correction [DH90a, DH90b, DH95, HSS99a]. deficient [Hem82e]. dependent [KH91a]. Derivation [HHL65]. derivative [Hem83a, Hem84d]. design [ELE^{+}06]. development [HKWdZ83]. device [Hem88a, Hem88b, Hem90b]. devoted [WHO08]. difference [HSS01b, HS01]. differences [Hem74a]. differentiaalvergelijkingen [BDHv75, DHv72]. differential [Hem71a, Hem72a, Hem72b]. Diffusion [HSS97b, Hem82d, HSS99a, HSS01b, HSS02a, HSS02c, HST02, HSS03a, HSS03b, KSS04, SSH04, VKH06, vRH05]. diffusional [HDH05]. Dimensional [Hd93a, Hd93b, HdZ97, Hem88a, Hem88b, Hem90b, HK91, HvR04, KdH88, KH91c]. Dimensions [KHd97]. diode [Hem90b]. direct [HSS03b]. direct-correction [HSS03b]. direction [KH91a]. directional [Hem82a]. Directions [KHd97]. Discontinuous [HS92, HS93, HS94a, HHvR03a, HHvR03b, HvR04, HHvR04, vRH05]. Discrete [FHS95, FHS96a, FHS96b, FHS96c, HS92, HS94a, HSS97b]. discretisation [HvR04]. Discretiseringsmethoden [BHPv76]. discretization [HHvR03a, HHvR03b, HHvR04]. discretized [vRH05]. distribution [FHST95]. disturbance [SSH97]. disturbances [HH09, HH13]. domain [FHST95, Hem74b, HST02]. double [HHv^{+}73]. double-length [HHv^{+}73].

During [HDH05]. dyadic [HS01]. dynamical [EHS95]. Dynamics [HKL^{+}95, HKL^{+}97a, HKL^{+}97b, DHKL94].

Editorial [WHO08]. Efficient [HK91, KH91b, KH92a, HS86, HKS86, KH91c]. elementary [HHv^{+}73]. elements [Hem73c, HHvR04]. ellipsometry [RTH^{+}11]. elliptic [Hem73c, Hem82a, Hem95a, KSS04]. embedded [HHvR03a]. EMG [WHO08]. enzymatic [HH09, HH13]. enzyme [HH69, HH09, HH13]. Enzymes [HHL65, HH68]. equation [Hem73c, Hem82d, HL08, KSS04, vRH05]. equations [DHK93, Hem71a, Hem72a, Hem72b, HS81, HS84, HS85a, HS85b, Hem86b, HS86, HKS86, HJ87, Hem87, HK88a, HK88b, HK88c, HK89, HM91, HK91, HD93c, HK94, HK95a, HK95b, HSS99b, HST02, MH90, vdMHKM93]. ERCIM [DHKL94]. Error [dH79]. errors [KvBH^{+}03]. estimation [vH75, Hem72a, Hem72b]. EULER [vdMHE90, HS84, HS85a, HS85b, Hem86a, Hem86b, HS86, HKS86, HJ87, Hem87, HK88a, HK88b, HK89, HK91, HK94, HK95a, HK95b, KH91c, KH92b, KHd97, vdMHKM93]. Euler-Flow [KHd97]. European [WHO08]. evaluation [HHv^{+}71]. exact [HL08]. exchange [TH91, TH94]. exponentially [dH79]. Extensions [Hem82c]. exterior [Hem74b]. external [HHA09].

fast [HKWdZ83]. Finite [Hem97, VH95, Hem73c, Hem94, Hem95b, Hem96a, HS01]. Finite-Volume [VH95, Hem94, Hem95b, Hem96a]. fitted [dH79]. Fix [KvBH^{+}03]. floating [SSH97]. Flow [KHd97, Hem86a, KdH88, KH91a, KH91c, KH92b, KvBH^{+}03]. Flows [VH95, DHK93, KH91b, KH92a]. fluid [DHKL94, KvBH^{+}03]. fluorescent [HHA09]. Formation [HEH^{+}67]. Formulas [HHL65]. Fourier [Hem80a, HHvR03b, HvR04, HHvR04]. Fouten [Hem90a]. Fredholm [HS81]. free [RTH^{+}11]. functions [HHv^{+}73, HS01].

Galerkin [dH79, Hem75, HHvR03a, HHvR03b, HvR04, HHvR04, vRH05]. Gas [HKL^{+}95, HKL^{+}97a, HKL^{+}97b, Hem86a]. Gauss [Hem82f, KdH88]. General [HH69]. generation [HDH05, HHA09, HHH72]. Godunov [KH97]. Godunov-type [KH97]. Grid [DHH^{+}91, Hem00, HHH00, Hem80b, Hem80c, Hem81b, HS81, Hem86b, HS86, HKS86, Hem94, Hem95b, NH97, NH98, NH00, Hem96a]. gridfunctions [Hem80a]. Grids [Hd93a, Hd93b, HdZ97, HKN97, KH97, HP96, HP97, HKN98, HS01].

High [HSS97a, HSS01a, HSS02c, HSS02b, HSS03a, KSS04, HSS99b, HSS00, HSS01b, HSS02a, HSS03b, RTH^{+}11]. high-affinity [RTH^{+}11]. High-Order [HSS97a, HSS01a, HSS02c, HSS02b, HSS03a, KSS04, HSS99b, HSS00, HSS01b, HSS02a, HSS03b]. higher [Hem86b, HvR04]. highest [Hem83a, Hem84d]. homogeneous [HHH72]. hoogleraar [Hem90a]. hyperbolic [DH90a, DH90b, DH95, HD93c]. hypersonic [KH91a, KH91b, KH92a].

Identification [HK93, EHS95]. II [FHS96b]. III [FHS96c]. ILU [Hem82f]. impacts [HH09, HH13]. implementations [HdZ85]. implicit [DH90a, DH90b]. improved [Hem87]. incomplete [Hem80b]. Index [Hem73a]. Industriële [Hem90a]. infarction [DHH^{+}02]. inhibition [HH68]. inhibitor [DHH^{+}02]. inhomogeneous [HP93, HHH72]. Initial [HS92, HS93, HS94a, vH75, FHST95, SSH97]. integral [HS81]. Interaction [HHL65, HH68]. interfaces [KvBH^{+}03]. interior [SSH04]. Introduction [Hem81a]. inverse [Hem82e]. ion [TH91, TH94]. ion-exchange [TH91, TH94]. issue [WHO08]. iteration [DH95, Hem82d, Hem84a]. iterative [DH90a, DH90b]. IV [HH68, HW94b].

John [HOS03].

kernel [HL08]. kind [HS81]. Kinetic

[HHL65, HH68]. Kinetics

[HK93, HH69, HH68]. KWIC [Hem73a].

Label [RTH^{+}11]. Label-free [RTH^{+}11]. layer [SSH04]. Layers [FHS95, FHS96a, FHS96b, FHS96c, Hem86a]. Lecture [Hem81b]. length [HHv^{+}73]. level [EH08, Hem80d, HHvR03b, HvR04, HHvR04, vRH05]. libraries [Hem77a]. library [Hem73a]. limits [WHH06]. Line [KdH88, Hem82f]. Line-Gauss [Hem82f]. Lineaire [Hem71b]. linear [ELE^{+}06, HdZ85, HK89, HHvR04]. Lobatto [Hem75]. local [FHST95]. LU [Hem80b]. LU-decomposition [Hem80b].

Maken [Hem90a]. Manifold [ELE^{+}06, EH08, HE07b, HE07a]. Manifold-mapping [ELE^{+}06, HE07a]. Manual [EHS95]. mapping [EH05, ELE^{+}06, EH08, ELH08a, ELH08b, HE07b, HE07a]. Math. [Hem96a]. Mathematical [HdH06, HKT93, HHA09]. measurement [HH09, HH13]. measuring [HHA09]. meerstapsmethoden [Hem71b]. Mesh [HS94b, PWBH81, SSH04]. Mesh-Parameter [PWBH81]. meshes [SSH97]. method [FHST95, Hem74a, Hem75, Hem80b, Hem82a, HKS86, Hem88b, HK89, Hem90b, HSS98, HSS99a, KSS04, PWBH81, SSH97, vRH05, vdMHKM93]. Methods [HSS97b, dH79, Hem72a, Hem81a, Hem81b, HS81, HKWdZ83, Hem83a, Hem84d, HW94b, HST02, TH91, TH94]. Miller [HOS03]. Mixed [Hem82d, Hem84a]. mixtures [HHH72]. Model [Hem00, HHH00, Hem96c, NH00]. models [HdH06]. modules [HWD84]. moving [SSH04]. MR1338896 [Hem96a]. Multi [Hem80c, KH91c, KH92b, Hem80b, Hem80d, Hem86b, HKS86, HK91]. Multi- [KH92b]. Multi-dimensional [KH91c, HK91]. Multi-grid [Hem80c, Hem80b, HKS86]. multi-level [Hem80d]. Multigrid [DHH^{+}91, HKWdZ83, Hem83a, Hem84b, Hem84c, Hem84d, HS84, HS85a, HS85b, HJ87, Hem87, HK88c, HW94a, HW94b, HKL^{+}95, HKN97, HKL^{+}97a, HKL^{+}97b, HK97, KHE97, VH95, WHO08, Hem81a, Hem82f, HWD84, HdZ85, HK88a, HK88b, Hem88a, Hem88b, HK89, HvdME90, Hem90b, Hem90c, HM91, Hem92, HK94, Hem94, HK95a, HK95b, Hem95b, Hem96a, Hem97, HKN98, HHvR03b, KdH88, KH91a, KH91b, KH91c, KH92a, MH90, VKH06, vRH05, vdMHKM93]. multigridding [KH92b]. multilevel [HE07b]. Multiple [HS81, HS86, KHE97, Hem81b]. myocardial [DHH^{+}02].

n [TH91]. n-component [TH91]. Navier [DHH^{+}91, DHK93, HK88c]. near [KvBH^{+}03]. Necessary [HS94b]. nested [Hem73b]. Neumann [HSS99b]. nieuw [Hem92]. no [Hem96a]. non [HK89]. non-linear [HK89]. Nonlinear [vH75, Hem90b, Hem72b, Hem73c, HKS86, HK88a, HK88b, Hem88b, HK94, HK95a, HK95b]. note [Hem82e]. notes [Hem81b]. Novel [HSS03b]. NUMAL [Hem73a, Hem81c]. Numerical [Hem72a, Hem77b, HM79, Hem81c, Hem83b, HSS97b, Hem73a, Hem74b, HW79, Hem82a, Hem96c, HSS99b, HST02, SSH97].

obtain [HHA09]. ODE [Hem83c]. one [Hem74a, Hem88a, Hem88b, Hem90b]. one-dimensional [Hem88a, Hem88b, Hem90b]. one-sided [Hem74a]. operator [Hem73c]. operators [HW79, HS01]. Oplossen [HR76]. optimization [ELE^{+}06, EH08, HE07b, HE07a]. Order [HSS97a, Hem86b, Hem90c, HSS99b, HSS00, HSS01b, HSS01a, HSS02a, HSS02c, HSS02b, HSS03a, HSS03b, KSS04]. Ordered [HKN97, HP96, HP97, HKN98]. Osher [HS86]. over-set [HHH00].

package [EHS95]. Parabolic [FHS95, FHS96a, HS92, HS93, HS94a, HSS97a, FHS96b, FHS96c, HS95, HSS97c, HSS98, HSS99b, HSS99a, HSS00, HSS01a, HSS02a, HSS02c, HSS02b, HSS03a, HSS03b]. parallel [HSS01b, HSS01a]. Parameter [Hem72b, HK93, HST02, PWBH81, vH75, EHS95, Hem72a, Hem83a, Hem84d]. Parameter-uniform [HST02]. Partially [HKN97, HP96, HP97, HKN98]. patients [DHH^{+}02]. PDEs [HS92, HS93, HS94a]. PEIA [RTH^{+}11]. PEIA-ellipsometry [RTH^{+}11]. Perturbation [HSS97a, Hem00, HHH00, Hem74b, HM79, Hem82a, HD83, Hem83b, Hem84a, HS95, HSS97c, HSS98, HSS00, HSS02b, HSS03a, NH00]. Perturbed [FHS95, FHS96a, HS92, HS94b, HS94a, HSS97b, FHST95, FHS96b, FHS96c, Hem83c, Hem96c, HSS99b, HSS01b, HSS01a, HSS02a, HSS02c, HST02, HSS03b, KSS04, SSH97, SSH04]. Pharmacokinetics [DHH^{+}02]. Pocklington [HL08]. Point [Hem77b, dH79, HSd80]. points [Hem74a, Hem75]. portable [HWD84]. posteriori [SSH04]. preliminaries [HKT93]. preliminary [HW79]. Principle [Hem82b, Hem82c]. Problem [Hem00, HHH00, Hem74b, Hem82a, HD83, Hem84a, Hem96c, HSS99b, HSS99a, NH00]. problemen [Hem92]. Problems [FHS95, FHS96a, Hem77b, HKL^{+}95, HSS97b, HKL^{+}97a, HKL^{+}97b, HSS97a, DH90a, DH90b, DH95, vH75, FHST95, FHS96b, FHS96c, dH79, Hem74a, HM79, HSd80, Hem83a, Hem83b, Hem84d, HS94b, Hem95b, HS95, Hem96a, Hem97, HSS97c, HSS98, HSS00, HSS01b, HSS01a, HSS02a, HSS02c, HSS02b, HSS03a, HSS03b, SSH97, SSH04, VKH06]. procedure [Hem71a, HHA09]. Procedures [Hem81c, CHvS72, Hem73a, Hem90c]. processes [Hem82e]. project [HK93]. prolongations [Hem80a, Hem90c]. protein [DHH^{+}02]. prothrombin [HEH^{+}67, HHv^{+}71].

quasilinear [HST02].

randwaardeproblemen [CHvS72, HR76]. rank [Hem82e]. Reaction [HK93, HH09, HH13, KSS04]. reaction-diffusion [KSS04]. rebuttal [HdH06]. rectangle [HSS98]. Rede [Hem90a]. refinement [Hem95a]. region [HE07a]. regular [HP96, HP97]. Rekenen [Hem96b]. relating [HH09, HH13]. relaxation [Hem80b, Hem82f, KdH88]. Remarks [Hem94]. report [HW79]. representation [HS01]. Research [DHKL94]. restrictions [Hem80a, Hem90c]. results [Hem90b, NH97, NH98]. Richardson [KSS04]. Robin [HSS02c, HSS03a]. rules [Hem73b]. Run [Hem84c, Hem84b].

scheme [HS86]. Schemes [DHH^{+}91, HSS97a, KH97, Hem86b, HK88c, HSS00, HSS01b, HSS01a, HSS02a, HSS02c, HSS02b, HSS03a, HSS03b, KH91c]. second [HS81]. Seidel [Hem82f, KdH88]. Semi [KHE97, KHd97, Hem95a]. Semi-Coarsened [KHE97]. Semi-Coarsening [KHd97]. semi-refinement [Hem95a]. semiconductor [Hem88a, Hem88b, Hem90b, HM91, MH90]. seminar [Hem81b]. sequence [Hem73b]. set [HHH00]. Sets [HKN97, HP96, HP97, HKN98]. shifts [HDH05]. sided [Hem74a]. simulation [Hem72a, Hem88a, Hem88b, Hem90b, WHH06]. Single [HHv^{+}73]. Single- [HHv^{+}73]. Singular [HSS97a, Hem00, HHH00, Hem74b, HM79, Hem82a, HD83, Hem83b, Hem84a, HS95, HSS97c, HSS98, HSS00, HSS02b, HSS03a, NH00]. Singularly [FHS95, FHS96a, HS92, HS94b, HS94a, HSS97b, FHST95, FHS96b, FHS96c, Hem83c, Hem96c, HSS99b, HSS01b, HSS01a, HSS02a, HSS02c, HST02, HSS03b, KSS04, SSH97, SSH04]. small [Hem83a, Hem84d]. software [EHS95]. solute [TH91, TH94]. Solution [DHH^{+}91, Hem95a, HSS97b, HKL^{+}97b, DHK93, Hem71a, Hem73c, Hem74b, HS81, Hem82a, Hem82d, HD83, Hem83c, Hem84a, HS85a, HS85b, Hem86b, HS86, HKS86, HM91, HS95, HSS97c, HSS99b, HL08, KvBH^{+}03, MH90, vdMHKM93]. Solution-Adaptive [HKL^{+}97b, vdMHKM93]. solutions [HS84]. Solvers [DHH^{+}91, HKWdZ83, HdZ85]. solving [HSS01b]. Some [HdZ85, HSd80]. soort [Hem92]. Space [EH05, ELH08a, ELH08b, HSS03b]. Sparse [Hd93a, Hd93b, Hem95b, HdZ97, Hem00, KH97, Hem94, NH97, NH98, NH00, Hem96a]. Sparse-Grid [Hem00, Hem95b, Hem94, NH00, Hem96a]. Special [WHO08]. spIds [EHS95]. SPR [RTH^{+}11]. stable [HHH72]. stage [HHv^{+}71]. Standard [HK97]. staplengte [Hem71b]. states [HHH72]. Steady [HKL^{+}95, HKL^{+}97a, HKL^{+}97b, HS84, HS85a, HS85b, Hem86b, HS86, HKS86, HK88a, HK88b, HK88c, HK89, HK91, HK94, HK95a, HK95b, KdH88, KH91b, KH92a, KH92b]. Stiff [Hem77b, dH79, Hem71a, Hem74a]. stijve [BDHv75, DHv72]. stimulated [DHKL94]. Stokes [DHH^{+}91, DHK93, HK88c]. strategy [HE07a, HL08, VKH06].

Structure [Hd93a, Hd93b, HdZ97, Hem80d, HvdME90]. Structured [VH95]. Study [Hem77b]. substrate [HH09, HH13]. substrates [HHA09]. Supercomputers [Hem84c, Hem84b]. system [Hem72a, HdZ85, WHH06]. systems [EHS95, Hem95a].

Technique [HS94b, Hem00, EH08, HHH00, HHA09, NH00, SSH04]. Techniques [HK97]. tests [HH68]. theory [HHH72]. Three [KHd97]. thrombin [HDH05, HHA09]. Time [HSS97a, HH09, HH13, HSS99b, HSS00, HSS01b, HSS01a, HSS02c, HSS02b, HSS03a, HSS03b]. Time-Accuracy [HSS97a, HSS00, HSS02b]. time-accurate [HSS01b, HSS01a, HSS02c, HSS03a]. Time-course [HH09, HH13]. transform [HHA09]. transport [TH91, TH94]. transportable [Hem77a]. trends [DHK93]. trust [HE07a]. trust-region [HE07a]. turning [Hem74a]. tweepunts [HR76]. Two [Hem77b, HHvR03b, vRH05, EH08, dH79, HHv^{+}71, HSd80, HvR04, HHvR04, KdH88, KvBH^{+}03]. two-dimensional [KdH88]. two-fluid [KvBH^{+}03]. Two-level [HHvR03b, vRH05, EH08, HvR04, HHvR04]. two-point [dH79, HSd80]. two-stage [HHv^{+}71]. type [KH97].

uitgesproken [Hem90a]. Uniform [HSS97a, HSS99a, HSS00, HST02]. Universiteit [Hem90a, Hem96b]. Upwind [DHH^{+}91, HK88c, KH91c]. upwinding [HK91, KH92b]. use [Hem83c, HS95, HSS97c].

Value [FHS95, FHS96a, Hem77b, HS94b, vH75, FHST95, FHS96b, FHS96c, dH79, Hem74a, HSd80]. variabele [Hem71b]. vector [HWD84]. vector-code [HWD84]. Verbeteren [Hem90a]. Volume [VH95, Hem94, Hem95b, Hem96a, Hem97]. voor [CHvS72, Hem92].

Wavelet [HP93]. Wavelets [HKT93]. weighted [Hem74a]. which [HS94b]. wire [HL08]. Wiskunde [Hem90a]. without [Hem82a, HHA09].

References

Beentjes:1975:CSD

[BDHv75] P. A. Beentjes, K. Dekker, P. W. Hemker, and M. van Veldhuizen. Colloquium stijve differentiaalvergelijkingen, deel 3, volume 15.3 of Mathematical Centre Syllabus. Mathematical Centre, Amsterdam, The Netherlands, 1975.

Bakker:1976:CD

[BHPv76] M. Bakker, P. W. Hemker, S. J. Polak, and M. van Veldhuizen. Colloquium Discretiseringsmethoden, volume 27 of Mathematical Centre Syllabus. Mathematical Centre, Amsterdam, The Netherlands, 1976.

Boehmer:1984:DCAa

[BHS84] K. Böhmer, P. W. Hemker, and H. J. Stetter. The defect correction approach. In K. Böhmer and H. J. Stetter, editors, Defect Correction Methods: Theory and Applications, volume 5 of Computing. Supplementum, pages 1–32. Springer-Verlag, Berlin, Heidelberg, New York, Tokyo, 1984. CODEN COSPDM. ISSN 0344-8029.

Coolen:1972:APV

[CHvS72] T. M. T. Coolen, P. W.

Hemker, P. J. van der Houwen, and E. Slagt. ALGOL 60 procedures voor begin- en randwaardeproblemen, volume 20 of Mathematical Centre Syllabus. Mathematical Centre, Amsterdam, The Netherlands, 1972.

Groen:1979:EBE

[dH79] P. P. N. de Groen and P. W.

Hemker. Error bouds for exponentially fitted Galerkin methods applied to stiff twopoint boundary-value problems. In P. W. Hemker and J. J. H. Miller, editors, Numerical Analysis of Singular Perturbation problems, pages 217– 249. Academic Press, London, 1979.

Desideri:1990:ACIa

[DH90a] J.-A. Désidéri and P. W.

Hemker. Analysis of the convergence of iterative implicit and defect-correction algorithms for hyperbolic problems. Technical Report NM-R9004, Centrum voor Wiskunde en Informatica, P. O. Box 4079, 1009 AB Amsterdam, The Netherlands, 1990.

Desideri:1990:ACIb

[DH90b] J.-A. Désidéri and P. W.

Hemker. Analysis of the convergence of iterative implicit and defect-correction algorithms for hyperbolic problems. Technical Report RR No 1200, INRIA, 1990.

Desideri:1995:CAD

[DH95] J.-A. Désidéri and P. W. Hemker. Convergence analysis of the defect-correction iteration for hyperbolic problems. SIAM Journal on Scientific Computing, 16(1):88–118, 1995. CODEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

Deconinck:1991:SAN

[DHH^{+}91] H. Deconinck, P. W. Hemker, Ch. Hirsch, M. Napolitano, and P. Larsen. Solution adaptive Navier–Stokes solvers with grid decoupled upwind schemes and multigrid acceleration. Technical report, Von Karman Institute, Rhode St. Genese, 1991.

Diris:2002:PCI

[DHH^{+}02] J. H. Diris, W. T. Hermens, P. W. Hemker, W. K. Lagrand, C. E. Hack, and M. P. van Dieijen-Visser. Pharmacokinetics of C1-inhibitor protein in patients with acute myocardial infarction. Clin Pharmacol Ther., 72(5):498–504, November 2002.

Deconinck:1993:NTS

[DHK93] H. Deconinck, P. W. Hemker, and B. Koren. New trends in the solution of the Navier– Stokes equations for compressible flows. In H. J. Allgeier, editor, Proceedings of the Aerodays 93, 2nd European Communities Aeronautics RTD Conference, Naples, 1993, pages 109–116. Commission of the European Communities, Luxembourg, 1993.

Desideri:1994:RCF

[DHKL94] Jean-Antoine Désidéri, Pieter W.

Hemker, Barry Koren, and Marie-Hélène Lallemand. Research in computational fluid dynamics, stimulated by ERCIM. In From universal morphisms to megabytes: a Baayen space odyssey, pages 269–286. Stichting Mathematisch Centrum, Centrum voor Wiskunde en Informatica, Amsterdam, The Netherlands, 1994.

Dekker:1972:CSD

[DHv72] Th.J. Dekker, P. W. Hemker, and P. J. van der Houwen. Colloquium stijve differentiaalvergelijkingen, deel 1, volume 15.1 of Mathematical Centre Syllabus. Mathematical Centre, Amsterdam, The Netherlands, 1972.

Echeverria:2005:SMDb

[EH05] D. Echeverría and P. W.

Hemker. Space mapping and defect correction. Computational Methods in Applied Mathematics, 5(2):107– 136 (electronic), 2005. ISSN 1609-4840 (print), 1609-9389 (electronic).

Echeverria:2008:MMT

[EH08] D. Echeverría and P. W.

Hemker. Manifold mapping: a two-level optimization technique. Computing and Visualization in Science, 11(4–6): 193–206, 2008. ISSN 1432-9360 (print), 1433-0369 (electronic).

Everaars:1995:MSS

[EHS95] C. T. H. Everaars, P. W. Hemker, and W. Stortelder. Manual of spIds, A software package for parameter identification in dynamical systems. Technical Report NM-R9521, Centrum voor Wiskunde en Informatica, P. O. Box 4079, 1009 AB Amsterdam, The Netherlands, 1995.

Echeverria:2006:MMO

[ELE^{+}06] D. Echeverria, D. Lahaye, L. Encica, E. A. Lomonova, P. W. Hemker, and A. J. A. Vandenput. Manifold-mapping optimization applied to linear actuator design. IEEE Transactions on Magnetics, 42(4): 1183–1186, April 2006. CO-DEN IEMGAQ. ISSN 0018-9464 (print), 1941-0069 (electronic).

Echeverria:2008:SMD

[ELH08a] David Echeverría, Domenico Lahaye, and Piet W. Hemker. Space mapping and defect correction. In Model order reduction: theory, research aspects and applications, volume 13 of Math. Ind., pages 157–176. Springer-Verlag, Berlin, Germany /Heidelberg, Germany / London, UK / etc., 2008.

Echeverria:2008:SMDa

[ELH08b] David Echeverría, Domenico Lahaye, and Piet W. Hemker. Space mapping and defect correction. In Wilhelmus H. A. Schilders et al., editors, Model order reduction: Theory, research aspects and applications. Selected papers based on the presentations at the workshop ‘Model order reduction, coupled problems and optimization’, Leiden, The Netherlands, September 19–23, 2005, volume 13 of Mathematics in Industry, pages 157–176. Springer-Verlag, Berlin, Germany /Heidelberg, Germany / London, UK / etc., 2008.

Farrell:1995:DAS

[FHS95] P. A. Farrell, P. W. Hemker, and G. I. Shishkin. Discrete approximations for singularly perturbed boundary value problems with parabolic layers. Technical Report NM-R9502, Centrum voor Wiskunde en Informatica, P. O. Box 4079, 1009 AB Amsterdam, The Netherlands, 1995.

Farrell:1996:DASa

[FHS96a] P. A. Farrell, P. W. Hemker, and G. I. Shishkin. Discrete approximations for singularly perturbed boundary value problems with parabolic layers, I. Journal of Computational Mathematics, 14(1):71– 97, 1996. CODEN JCMMEB. ISSN 0254-9409 (print), 1991-7139 (electronic).

Farrell:1996:DASb

[FHS96b] P. A. Farrell, P. W. Hemker, and G. I. Shishkin. Discrete approximations for singularly perturbed boundary value problems with parabolic layers. II. Journal of Computational Mathematics, 14(2):183– 194, 1996. CODEN JCMMEB. ISSN 0254-9409 (print), 1991-7139 (electronic).

Farrell:1996:DASc

[FHS96c] P. A. Farrell, P. W. Hemker, and G. I. Shishkin. Discrete approximations for singularly perturbed boundary value problems with parabolic layers. III. Journal of Computational Mathematics, 14(3):273– 290, 1996. CODEN JCMMEB. ISSN 0254-9409 (print), 1991-7139 (electronic).

Farrell:1995:DDM

[FHST95] P. A. Farrell, P. W. Hemker, G. I. Shishkin, and I. V. Tselishcheva. A domain decomposition method for singularly perturbed boundary value problems with a local distribution of the initial conditions. In Abstracts of the International Conference on Advanced Mathematics, V. Computations and Applications, page 102. ????, ????, 1995. Novosibirsk, June 20–24.

Hemker:1983:DCSa

[HD83] P. W. Hemker and P. M. De Zeeuw. Defect correction for the solution of a singular perturbation problem. In R. S. Stepleman, editor, Scientific Computing, pages 113– 118. North-Holland, Amsterdam, The Netherlands, 1983.

Hemker:1993:BDSa

[Hd93a] P. W. Hemker and P. M. de Zeeuw. BASIS3: A data structure for 3-dimensional sparse grids. Technical Report CR-1993-30, Von Karman Institute, Rhode St. Genese, 1993.

Hemker:1993:BDSb

[Hd93b] P. W. Hemker and P. M. de Zeeuw. BASIS3: A data structure for 3-dimensional sparse grids. Technical Report NM-R9321, Centrum voor Wiskunde en Informatica, P. O. Box 4079, 1009 AB Amsterdam, The Netherlands, 1993.

Hemker:1993:CBD

[HD93c] P. W. Hemker and J.-A.

Désidéri. Convergence behaviour of defect correction for hyperbolic equations. Journal of Computational and Applied Mathematics, 45(3):357– 365, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Hemker:2005:DCT

[HDH05] H. C. Hemker, E. De Smedt, and P. W. Hemker. During coagulation, thrombin generation shifts from chemical to diffusional control. Journal of Thrombosis and Haemostasis:

JTH, 3(11):2399–2400, November 2005. ISSN 1538-7836 (print), 1538-7933 (electronic).

Hemker:2006:MBM

[HdH06] H. C. Hemker, E. de Smedt, and P. W. Hemker. Mathematical and biological models of blood coagulation. a rebuttal. Journal of Thrombosis and Haemostasis: JTH, 4(3):710– 711, March 2006. ISSN 1538-7836 (print), 1538-7933 (electronic).

Hemker:1985:SIMa

[HdZ85] P. W. Hemker and P. M. de Zeeuw. Some implementations of multigrid linear system solvers. In D. J. Paddon and H. Holstein, editors, Multigrid methods for integral and differential equations (Bristol, 1983), volume 3 of The Institute of Mathematics and its Applications Conference Series, pages 85–116. Clarendon Press, Oxford, UK, 1985.

Hemker:1997:BDS

[HdZ97] P. W. Hemker and P. M. de Zeeuw. BASIS3: A data structure for 3-dimensional sparse grids. In Euler and Navier–Stokes Solvers Using Multi-Dimensional Upwind Schemes and Multigrid Acceleration, volume 57 of Notes on Numerical Fluid Mechanics, pages 443–484. Friedrich Vieweg und Sohn, Braunschweig, Germany, 1997.

Hemker:2007:TRS

[HE07a] P. W. Hemker and D. Echeverría. A trust-region strategy for manifold-mapping optimization. Journal of computational physics, 224(1):464–475, May 20, 2007. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0021999107001519.

Hemker:2007:MMM

[HE07b] Pieter W. Hemker and David Echeverría. Manifold mapping for multilevel optimization. In Scientific computing in electrical engineering, volume 11 of Math. Ind., pages 325–330. Springer-Verlag, Berlin, Germany /Heidelberg, Germany / London, UK / etc., 2007.

Hemker:1967:FPC

[HEH^{+}67] H. C. Hemker, M. P. Esnouf, P. W. Hemker, A. C. W. Swart, and R. G. MacFarlane. Formation of prothrombin converting activity. Nature, 215(5098): 248–251, July 15, 1967. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Hemker:1971:APS

[Hem71a] P. W. Hemker. An ALGOL 60 procedure for the solution of stiff differential equations. Technical Report MR 128, Mathematical Centre, Amsterdam, The Netherlands, 1971.

Hemker:1971:LMM

[Hem71b] P. W. Hemker. Lineaire meerstapsmethoden met variabele staplengte. Technical Report NR 15, Mathematical Centre, Amsterdam, The Netherlands, 1971.

Hemker:1972:NMD

[Hem72a] P. W. Hemker. Numerical methods for differential equations in system simulation and in parameter estimation. In H. C. Hemker and B. Hess, editors, Analysis and Simulation of biochemical systems, pages 59–80. North-Holland, Amsterdam, The Netherlands, 1972.

Hemker:1972:PEN

[Hem72b] P. W. Hemker. Parameter estimation in nonlinear differential equations. Technical Report MR 134, Mathematical Centre, Amsterdam, The Netherlands, 1972.

Hemker:1973:NLN

[Hem73a] P. W. Hemker. NUMAL, a library of numerical procedures in ALGOL 60: Index and KWIC index. Technical Report NW 8, Mathematical Centre, Amsterdam, The Netherlands, 1973.

Hemker:1973:SNC

[Hem73b] P. W. Hemker. A sequence of nested cubature rules. Technical Report NW 3, Mathematical Centre, Amsterdam, The Netherlands, 1973.

Hemker:1973:SNE

[Hem73c] P. W. Hemker. The solution of a nonlinear elliptic operator equation by finite elements. Technical Report NW 6, Mathematical Centre, Amsterdam, The Netherlands, 1973.

Hemker:1974:MWO

[Hem74a] P. W. Hemker. A method of weighted one-sided differences for stiff boundary value problems with turning points. Technical Report NW 9, Mathematical Centre, Amsterdam, The Netherlands, 1974.

Hemker:1974:NSS

[Hem74b] P. W. Hemker. The numerical solution of a singular perturbation problem in the domain exterior of a circle. Technical Report NW 11, Mathematical Centre, Amsterdam, The Netherlands, 1974.

Hemker:1975:GML

[Hem75] P. W. Hemker. Galerkin’s method and Lobatto points. Technical Report NW 24, Mathematical Centre, Amsterdam, The Netherlands, 1975.

Hemker:1977:CTA

[Hem77a]  P. W. Hemker.  Criteria

for transportable ALGOL libraries. In W. Cowell, editor, Portability of numerical software, volume 57 of Lecture Notes in Computer Science, pages 145–157. Springer-Verlag, Berlin, Germany /Heidelberg, Germany / London, UK / etc., 1977.

Hemker:1977:NSS

[Hem77b] P. W. Hemker. A Numerical Study of Stiff Two Point Boundary Value Problems, volume 80 of Mathematical Centre Tract. Mathematical Centre, Amsterdam, The Netherlands, 1977.

Hemker:1980:FAG

[Hem80a] P. W. Hemker. Fourier analysis of gridfunctions, prolongations and restrictions. Technical Report NW 93, Mathematical Centre, Amsterdam, The Netherlands, 1980.

Hemker:1980:ILDa

[Hem80b] P. W. Hemker. The incomplete LU-decomposition as a relaxation method in multi-grid algorithms. In J. J. H. Miller, editor, Boundary and Interior Layers—Computational and Asymptotic Methods, pages 306–311. Boole Press, Dublin, Ireland, 1980.

Hemker:1980:MGBa

[Hem80c] P. W. Hemker. Multi-grid bibliography. In J. Verwer, editor, Colloquium Numerical Integration of Partial Differential Equations. Dept. of Numerical Mathematics, Mathematical Center, Amsterdam, The Netherlands, 1980.

Hemker:1980:SAM

[Hem80d] P. W. Hemker. On the structure of an adaptive multi-level algorithm. BIT, 20(3):289– 301, 1980. CODEN BIT-TEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Hemker:1981:IMMa

[Hem81a] P. W. Hemker. Introduction to multigrid methods. Nieuw Archief voor Wiskunke, 3:71– 101, 1981. CODEN NAWIA7. ISSN 0028-9825.

Hemker:1981:LNS

[Hem81b] P. W. Hemker. Lecture notes of a seminar on multiple grid methods. Technical Report NN 24, Mathematical Centre, Amsterdam, The Netherlands, 1981.

Hemker:1981:NNP

[Hem81c] P. W. Hemker. NUMAL, Numerical Procedures in ALGOL 60, volume 47.1–47.7 of Mathematical Centre Syllabus. Mathematical Centre, Amsterdam, The Netherlands, 1981. 7 Vols.:

1. Elementary procedures, 2. Algebraic evaluations, 3. Linear Algebra, 4. Analytic Evaluations, 5. Analytic problems,

6. Special Functions, 7. Interpolation and Approximation.

Hemker:1982:AMD

[Hem82a] P. W. Hemker. An accurate method without directional bias for the numerical solution of a 2-D elliptic singular perturbation problem. Lecture Notes in Mathematics, 942:192–206, 1982. CODEN LNMAA2. ISBN 3-540-11584-6 (print), 3-540-39332-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.
springer.com/chapter/10.
1007/BFb0094748/.

Hemker:1982:DCP

[Hem82b] P. W. Hemker. The defect correction principle. In J. J. H. Miller, editor, An Introduction to Computational and Asymptotic Methods for Boundary And Interior Layers, pages 11– 45. Boole Press, Dublin, 1982.

Hemker:1982:EDCa

[Hem82c] P. W. Hemker. Extensions of the defect correction principle. In J. J. H. Miller, editor, An Introduction to Computational and Asymptotic Methods for Boundary and Interior Layers, Short Course, pages 33–45. Boole Press, Dublin, Ireland, 1982.

Hemker:1982:MDC

[Hem82d] P. W. Hemker. Mixed defect correction iteration for the accurate solution of the convection diffusion equation. Lecture Notes in Mathematics, 960:485–501, 1982. CO-DEN LNMAA2. ISBN 3-540-11955-8 (print), 3-540-39544-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.

springer.com/chapter/10.
1007/BFb0069940/.

Hemker:1982:NDCa

[Hem82e] P. W. Hemker. A note on defect correction processes with an approximate inverse of deficient rank. Applied Mathematics and Computation, 8:137– 139, 1982. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Hemker:1982:CLGa

[Hem82f] P. W. Hemker. On the comparison of Line-Gauss–Seidel and ILU relaxation in multigrid algorithms. In J. J. H. Miller, editor, Computational and Asymptotic Methods for Boundary and Interior Layers, pages 269–277. Boole Press, Dublin, Ireland, 1982.

Hemker:1983:MMPa

[Hem83a] P. W. Hemker. Multigrid methods for problems with a small parameter in the highest derivative. In G. A. Watson, editor, Numerical Analysis, volume 1066 of Lecture Notes in Mathematics, pages 106– 121. Springer-Verlag, Berlin, Heidelberg, New York, Tokyo, 1983.

Hemker:1983:NAS

[Hem83b]  P. W. Hemker.  Numeri-

cal aspects of singular perturbation problems. Lecture Notes in Mathematics, 985:267–287, 1983. CO-DEN LNMAA2. ISBN 3-540-12286-9 (print), 3-540-39612-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

                                        tronic). URL http://link.
                                        springer.com/chapter/10.
                                        1007/BFb0062371/.

Hemker:1983:UDCa

[Hem83c] P. W. Hemker. The use of defect correction for the solution of a singularly perturbed ODE. In R. März, editor, Proceedings of the Third Conference on the Numerical Treatment of Ordinary Differential Equations, pages 91–104. Humboldt Universität, Berlin, 1983.

Hemker:1984:MDCa

[Hem84a] P. W. Hemker. Mixed defect correction iteration for the solution of a singular perturbation problem. In K. Böhmer and H. J. Stetter, editors, Defect Correction Methods: Theory and Applications, volume 5 of Computing. Supplementum, pages 123– 145. Springer-Verlag, Berlin, Heidelberg, New York, Tokyo, 1984. CODEN COSPDM. ISSN 0344-8029.

Hemker:1984:MARa

[Hem84b] P. W. Hemker. Multigrid algorithms run on supercomputers. Supercomputer, 4:44–51, 1984. CODEN SPCOEL. ISSN 0168-7875.

Hemker:1984:MARb

[Hem84c] P. W. Hemker. Multigrid algorithms run on supercomputers. CWI-Newsletter, 3:25–30, 1984.

Hemker:1984:MMP

[Hem84d]  P. W. Hemker.  Multi-

grid methods for problems with a small parameter in the highest derivative. Lecture Notes in Mathematics, 1066:106–121, 1984. CO-DEN LNMAA2. ISBN 3-540-13344-5 (print), 3-540-38881-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.
springer.com/chapter/10.
1007/BFb0099520/.

Hemker:1986:CLE

[Hem86a] P. W. Hemker. Computation of layers in Euler gas flow. In J. J. H. Miller S. K. Godunov and V. A. Novikov, editors, BAIL IV, Proceedings of the Fourth International Conference on Boundary and Interior Layers, Novosibirsk., pages 320–324. Boole Press, Dublin, Ireland, 1986.

Hemker:1986:DCH

[Hem86b] P. W. Hemker. Defect correction and higher order schemes for the multi grid solution of the steady Euler equations. Lecture Notes in Mathematics, 1228:149–165, 1986. CODEN LNMAA2. ISBN 3-540-17198-3 (print), 3-540-47372-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.
springer.com/chapter/10.
1007/BFb0072646/.

Hemker:1987:MIAa

[Hem87] P. W. Hemker. Multigrid and improved accuracy for the Euler equations. In E. Adams, R. Ansorge, Ch. Grossmann, and H.-G. Roos, editors, Discretization in Differential Equations and Enclosures (Weissig, 1986), volume 36 of Math. Res., pages 123–136. Akademie-Verlag, Berlin, Germany, 1987.

Hemker:1988:MAOa

[Hem88a] P. W. Hemker. A multigrid approach for one-dimensional semiconductor device simulation. In W. Joppich and U. Trottenberg, editors, Device Simulation –Numerische Problemfelder–, volume 144 of GMD-Studien. GMD, Bonn, Germany, 1988.

Hemker:1988:NMMa

[Hem88b] P. W. Hemker. A nonlinear multigrid method for onedimensional semiconductor device simulation. In Guo Ben-Yu, J. J. H. Miller, and Shi Zhong-ci, editors, BAIL V, Proceedings of the Fifth International Conference on Boundary and Interior Layers, Shanghai, 1988, volume 12 of Boole Press Conf. Ser., pages 18–29. Boole Press, Dublin, Ireland, 1988.

Hemker:1990:FMV

[Hem90a] P. W. Hemker. Fouten Maken en Verbeteren. Rede uitgesproken bij de aanvaarding van het ambt van Bijzonder hoogleraar in de Industriële Wiskunde aan de Universiteit van Amsterdam, Juni 12, 1990.

Hemker:1990:NMMa

[Hem90b] P. W. Hemker. Nonlinear multigrid method for onedimensional semiconductor device simulation: results for the diode. Journal of Computational and Applied Mathematics, 30:117–126, 1990. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Hemker:1990:OPRa

[Hem90c] P. W. Hemker. On the order of prolongations and restrictions in multigrid procedures. Journal of Computational and Applied Mathematics, 32(3):423– 429, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Hemker:1992:NSM

[Hem92] P. W. Hemker. Een nieuw soort multigrid voor problemen in 3D. In Is er nog nieuws? Universiteit van Amsterdam, Amsterdam, The Netherlands, 1992. Bijdrage in het Liber Amicorum, ter gelegenheid van het afscheid van prof. T.J. Dekker.

Hemker:1994:RSG

[Hem94]  P. W. Hemker.  Remarks

on sparse-grid finite-volume multigrid. Technical Report NM-R9421, Centrum voor Wiskunde en Informatica, P. O. Box 4079, 1009 AB Amsterdam, The Netherlands, 1994.

Hemker:1995:SES

[Hem95a] P. W. Hemker. Solution of 3D elliptic systems by semirefinement. Analele Stiintifice ale Universitatii Ovidius Constanta, Seria Matematica, 3(1): 63–82, 1995. ISSN 1223-723X.

Hemker:1995:SGF

[Hem95b] P. W. Hemker. Sparse-grid finite-volume multigrid for 3Dproblems. Advances in computational mathematics, 4(1–2): 83–110, December 1995. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF02123474. ???

Hemker:1996:CSG

[Hem96a] P. W. Hemker. Corrigendum: “Sparse-grid finitevolume multigrid for 3Dproblems” [Adv. Comput. Math. 4 (1995), no. 1–2, 83– 110; MR1338896 (96c:65193)]. Advances in computational mathematics, 5(1):125–126, December 1996. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF02124737;   http:/
/link.springer.com/content/
pdf/10.1007/BF02124737.pdf.

See [Hem95b].

Hemker:1996:RAU

[Hem96b] P. W. Hemker. Rekenen aan de universiteit. Nieuwsbrief van de vakgroep Wiskunde, september:5–7, 1996.

Hemker:1996:SPM

[Hem96c] P. W. Hemker. A singularly perturbed model problem for numerical computation. Journal of Computational and Applied Mathematics, 76 (1–2):277–285, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Hemker:1997:FVM

[Hem97]  P. W. Hemker.  Finite

volume multigrid for 3Dproblems. In Euler and Navier– Stokes Solvers Using Multi-Dimensional Upwind Schemes and Multigrid Acceleration, volume 57 of Notes on Numerical Fluid Mechanics, pages 485–509. Friedrich Vieweg und Sohn, Braunschweig, Germany, 1997.

Hemker:2000:AAS

[Hem00] P. W. Hemker. Application of an adaptive sparse-grid technique to a model singular perturbation problem. Computing: Archiv für informatik und numerik, 65(4):357–378, December 2000. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL

http://link.springer-ny.
com/link/service/journals/
00607/bibs/0065004/00650357.
htm;   http://link.springer-

ny.com/link/service/journals/
00607/papers/0065004/00650357.
pdf.

Hemker:1968:KAI

[HH68] H. C. Hemker and P. W. Hemker. Kinetic aspects of the interaction of blood clotting enzymes. IV. Kinetics of competitive inhibition in clotting tests. Thrombosis et diathesis haemorrhagica, 19(3):364–367, July 31, 1968. CODEN TD-HAAT. ISSN 0340-5338.

Hemker:1969:GKE

[HH69] H. C. Hemker and P. W. Hemker. General kinetics of enzyme cascades. Proceedings of the Royal Society of London. Series B. Biological sciences, 173(1032):411–420, July 1, 1969. CODEN PRLBA4. ISSN 0962-8452 (print), 1471-2954 (electronic).

Hemker:2009:TCM

[HH09] Hendrik Coenraad L. M. Hemker and Pieter Wilhelm Hemker. Time-course measurement of enzymatic activity corrected for impacts of disturbances relating to the reaction of the enzyme with a substrate. Official Gazette of the United States Patent and Trademark Office Patents, ??(??):??, August 13, 2009. CODEN OGUPE7. ISSN 0098-1133 (print), 1554-9550 (electronic). URL https:

//www.google.com/patents/
WO2009098313A1?cl=en. Patent

WO2009098313 A1, filed 7 February 2008.

Hemker:2013:TCM

[HH13] H. C. L. M. Hemker and P. W.

Hemker. Time-course measurement of enzymatic activity corrected for impacts of disturbances relating to the reaction of the enzyme with a substrate, October 8, 2013.

URL https://www.google.
com/patents/US8551722. US

Patent 8,551,722, filed 6 February 2009.

Hemker:2009:TMT

[HHA09] H. Coen Hemker, Pieter W.

Hemker, and Raed Al Dieri. The technique of measuring thrombin generation with fluorescent substrates: 4. The H-transform, a mathematical procedure to obtain thrombin concentrations without external calibration. Thrombosis and Haemostasis, 101(1):171– 177, January 2009. ISSN 0340-6245. Erratum in 101(2):412, February 2009.

Hermens:1972:TGS

[HHH72] W.Th. Hermens, H. C. Hemker, and P. W. Hemker. On the theory of the generation of stable inhomogeneous states in homogeneous mixtures. In H. C. Hemker and B. Hess, editors, Analysis and Simulation of biochemical systems, pages 19–37. North-Holland, Amsterdam, The Netherlands, 1972.

Havik:2000:ASG

[HHH00] E. D. Havik, P. W. Hemker, and W. Hoffmann. Application of the over-set grid technique to a model singular perturbation problem. Computing: Archiv für informatik und numerik, 65(4):339–356, 2000. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL

                                        http://link.springer-ny.
                                        com/link/service/journals/
                                        00607/bibs/0065004/00650339.
                                        htm;   http://link.springer-
                                        ny.com/link/service/journals/
                                        00607/papers/0065004/00650339.
                                        pdf.

Hemker:1965:KAI

[HHL65] H. C. Hemker, P. W. Hemker, and E. A. Loeliger. Kinetic aspects of the interaction of blood clotting enzymes. I. Derivation of basic formulas. Thrombosis et diathesis haemorrhagica, 13(??):155– 175, March 15, 1965. CODEN TDHAAT. ISSN 0340-5338.

Hemker:1971:ETS

[HHv^{+}71] H. C. Hemker, P. W. Hemker, K. van de Torren, P. P. Devillee, W. Th. Hermens, and E. A. Loeliger. The evaluation of a two-stage prothrombin assay. Thrombosis et diathesis haemorrhagica, 25(3):545–554, June 1971. CODEN TDHAAT. ISSN 0340-5338.

Hemker:1973:SDL

[HHv^{+}73] P. W. Hemker, W. Hoffmann, S. P. N. van Kampen, H. L. Oudshoorn, and D. T. Winter. Single- and double-length computation of elementary functions. Technical Report NW 7, Mathematical Centre, Amsterdam, The Netherlands, 1973.

Hemker:2003:DGD

[HHvR03a] P. W. Hemker, W. Hoffmann, and M. H. van Raalte. Discontinuous Galerkin discretization with embedded boundary conditions. Computational Methods in Applied Mathematics, 3 (1):135–158 (electronic), 2003. ISSN 1609-4840 (print), 1609-9389 (electronic). Dedicated to Raytcho Lazarov.

Hemker:2003:TLF

[HHvR03b] P. W. Hemker, W. Hoffmann, and M. H. van Raalte. Twolevel Fourier analysis of a multigrid approach for discontinuous Galerkin discretization. SIAM Journal on Scientific Computing, 25(3): 1018–1041 (electronic), 2003. CODEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/40510.

Hemker:2004:FTLb

[HHvR04] P. W. Hemker, W. Hoffmann, and M. H. van Raalte. Fourier two-level analysis for discontinuous Galerkin discretization with linear elements. Numerical Linear Algebra with Applications, 11(5–6):473–491, June/August 2004. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Hemker:1987:MAEa

[HJ87] P. W. Hemker and G. M. Johnson. Multigrid approaches to the Euler equations. In S. F. McCormick, editor, Multigrid Methods, volume 3 of Frontiers in Applied Mathematics, pages 57–72. SIAM, Philadelphia, PA, USA, 1987.

Hemker:1988:DCNa

[HK88a] P. W. Hemker and B. Koren. Defect correction and nonlinear multigrid for the steady Euler equations. Technical Report NM-N8801, Centrum voor Wiskunde en Informatica, P. O. Box 4079, 1009 AB Amsterdam, The Netherlands, 1988. Lecture Notes of a course given at VKI, March 7–11, 1988.

Hemker:1988:DCNb

[HK88b] P. W. Hemker and B. Koren. Defect correction and nonlinear multigrid for the steady Euler equations. In Computational fluid dynamics, Vol. 1, 2, volume 88 of von Karman Inst. Fluid Dynam. Lecture Ser., page 59. von Karman Inst. Fluid Dynamics, Rhode, 1988.

Hemker:1988:MDCa

[HK88c] P. W. Hemker and B. Koren. Multigrid, defect correction and upwind schemes for the steady Navier–Stokes equations. In K. W. Morton and M. J. Baines, editors, Numerical Methods for Fluid Dynamics, III (Oxford, 1988), volume 17 of Inst. Math. Appl. Conf. Ser. New Ser., pages 153–170. Clarendon Press, Oxford, UK, 1988.

Hemker:1989:NLM

[HK89] P. W. Hemker and B. Koren.

A non-linear multigrid method for the steady Euler equations. In A. Dervieux, B. van Leer, J. Périaux, and A. Rizzi, editors, Numerical Simulation of Compressible Euler Flows, volume 26 of Notes on Numerical Fluid Mechanics, pages 175–196. Friedrich Vieweg und Sohn, Braunschweig, Germany, 1989.

Hemker:1991:EMD

[HK91] P. W. Hemker and B. Koren.

Efficient multi-dimensional upwinding for the steady Euler equations. Technical Report NM-R9107, Centrum voor Wiskunde en Informatica, P. O. Box 4079, 1009 AB Amsterdam, The Netherlands, 1991.

Hemker:1993:PPI

[HK93] P. W. Hemker and J. Kok.

A project on parameter identification in reaction kinetics. Technical Report NM-R9301, Centrum voor Wiskunde en Informatica, P. O. Box 4079, 1009 AB Amsterdam, The Netherlands, 1993.

Hemker:1994:DCN

[HK94] P. W. Hemker and B. Koren. Defect correction and nonlinear multigrid for steady Euler equations. In W. G. Habashi, editor, Proceedings of the International Workshop on Solution Techniques for Large-Scale CFD Problems, Montréal, 1994, pages 223–241. Centre de Recherche en Calcul Appliqué, Montréal, 1994.

Hemker:1995:DCNa

[HK95a] P. W. Hemker and B. Koren. Defect correction and nonlinear multigrid for steady Euler equations. In W. G. Habashi, editor, Solution Techniques for Large-scale CFD Problems, pages 273–291. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, 1995.

Hemker:1995:DCNb

[HK95b] P. W. Hemker and B. Koren. Defect correction and nonlinear multigrid for steady Euler equations. In W. G. Habashi and M. M. Hafez, editors, Computational Fluid Dynamics Techniques, pages 699– 718. Gordon and Breach, Basel, 1995.

Hemker:1997:SMT

[HK97] P. W. Hemker and B. Koren. Standard multigrid techniques for CFD. In H. Deconinck, editor, Lecture Series on Computational Fluid Dynamics, volume 1997-02. von Karman Institute for Fluid Dynamics, Rhode Saint Genèse – Belgium, 1997.

Hemker:1995:MSG

[HKL^{+}95] P. W. Hemker, B. Koren, W. M. Lioen, M. Nool, and H. T. M. van der Maarel. Multigrid for steady gas dynamics problems. In M. Hafez and K. Oshima, editors, Computational Fluid Dynamics Review 1995, volume 1, pages 477–494. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, 1995. ISBN 0-471-95589-2. LCCN TA357 .C58785x.

Hemker:1997:MSG

[HKL^{+}97a] P. W. Hemker, B. Koren, W. M. Lioen, M. Nool, and H. T. M. van der Maarel. Multigrid for steady gas dynamics problems. In Euler and Navier–Stokes Solvers Using Multi-Dimensional Upwind Schemes and Multigrid Acceleration, volume 57 of Notes on Numerical Fluid Mechanics, pages 393–417. Friedrich Vieweg und Sohn, Braunschweig, Germany, 1997.

Hemker:1997:SAM

[HKL^{+}97b] P. W. Hemker, B. Koren, W. M. Lioen, M. Nool, and H. T. M. van der Maarel. Solution-adaptive multigrid for steady gas dynamics problems. In H. Deconinck, editor, Lecture Series on Computational Fluid Dynamics, volume 1997-02. von Karman Institute for Fluid Dynamics, Rhode Saint Genèse – Belgium, 1997.

Hemker:1997:MPO

[HKN97] P. W. Hemker, B. Koren, and J. Noordmans. 3D multigrid on partially ordered sets of grids. In H. Deconinck, editor, Lecture Series on Computational Fluid Dynamics, volume 1997-02. von Karman Institute for Fluid Dynamics, Rhode Saint Genèse – Belgium, 1997.

Hemker:1998:MPO

[HKN98] P. W. Hemker, B. Koren, and J. Noordmans. 3D multigrid on partially ordered sets of grids. In Multigrid methods V (Stuttgart, 1996), volume 3 of Lect. Notes Comput. Sci. Eng., pages 105–124. Springer-Verlag, Berlin, Germany /Heidelberg, Germany / London, UK / etc., 1998.

Hemker:1986:NMGa

[HKS86] P. W. Hemker, B. Koren, and S. P. Spekreijse. A nonlinear multi-grid method for the efficient solution of the steady Euler equations. In F. G. Zhuang and Y. L. Zhu, editors, Tenth International Conference on Numerical Methods in Fluid Dynamics, volume 264 of Lecture Notes in Physics, pages 308– 313. Springer-Verlag, Berlin, Heidelberg, New York, Tokyo, 1986.

Hemker:1993:WMP

[HKT93] Pieter W. Hemker, Tom H.

Koornwinder, and Nico M. Temme. Wavelets: mathematical preliminaries. In T. H. Koornwinder, editor, Wavelets: An Elementary Treatment of Theory and Applications, volume 1 of Ser. Approx. Decompos., pages 13–26. World Scientific Publishing Co., Singapore, Philadelphia, River Edge, NJ, 1993.

Hemker:1983:MMDa

[HKWdZ83] P. W. Hemker, R. Kettler, P. Wesseling, and P. M. de Zeeuw. Multigrid methods: development of fast solvers. Applied Mathematics and Computation, 13(3–4): 311–326, November 1983. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Hemker:2008:ASS

[HL08] P. W. Hemker and D. J. P. Lahaye. An hp-adaptive strategy for the solution of the exact kernel curved wire Pocklington equation. Computational Methods in Applied Mathematics, 8(1):39–59, 2008. ISSN 1609-4840 (print), 1609-9389 (electronic).

Hemker:1979:NAS

[HM79] P. W. Hemker and J. J. H.

Miller, editors. Numerical analysis of singular perturbation problems. Academic Press, New York, NY, USA, 1979.

ISBN 0-12-340250-6. LCCN QA371 .N79 1978.

Hemker:1991:AMAa

[HM91] P. W. Hemker and J. Molenaar. An adaptive multigrid approach for the solution of the 2D semiconductor equations. In W. Hackbusch and U. Trottenberg, editors, Multigrid methods, III (Bonn, 1990), volume 98 of International Series of Numerical Mathematics, pages 41–60. Birkhäuser Verlag, Basel, 1991.

Hemker:2003:JM

[HOS03] P. W. Hemker, E. O’Riordan, and G. I. Shishkin. John Miller — 65. Computational Methods in Applied Mathematics, 3 (3):359–360 (electronic), 2003. ISSN 1609-4840 (print), 1609-9389 (electronic). Dedicated to John J. H. Miller on the occasion of his 65th birthday.

Hemker:1993:WBA

[HP93] Pieter W. Hemker and Frédérique Plantevin. Wavelet bases adapted to inhomogeneous cases. In T. H. Koornwinder, editor, Wavelets: An Elementary Treatment of Theory and Applications, volume 1 of Ser. Approx. Decompos., pages 107– 128. World Scientific Publishing Co., Singapore, Philadelphia, River Edge, NJ, 1993.

Hemker:1996:APO

[HP96] P. W. Hemker and C. Pflaum. Approximation on partially ordered sets of regular grids. Technical Report NM-R9611, Centrum voor Wiskunde en Informatica, P. O. Box 4079, 1009 AB Amsterdam, The Netherlands, 1996.

Hemker:1997:APO

[HP97] P. W. Hemker and C. Pflaum.

Approximation on partially ordered sets of regular grids. Applied Numerical Mathematics: Transactions of IMACS, 25(1):55–87, 1997. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). to appear.

Hemker:1976:OVT

[HR76] P. W. Hemker and J. P.

Roos. Oplossen van tweepunts randwaardeproblemen. In J. Bus, editor, Colloquium Numerieke Programmatuur, volume 29 of Mathematical Centre Syllabus, pages 101–142. Mathematical Centre, Amsterdam, The Netherlands, 1976.

Hemker:1981:MGMa

[HS81] P. W. Hemker and H. Schippers. Multiple grid methods for the solution of Fredholm integral equations of the second kind. Mathematics of Computation, 36(153):215–232, January 1981. CODEN MCMPAF. ISSN 0025-5718 (paper), 1088-6842 (electronic).

Hemker:1984:MSSa

[HS84] P. W. Hemker and S. P.

Spekreijse. Multigrid solutions of the steady Euler equations. In D. Braess, W. Hackbusch, and U. Trottenberg, editors, Advances in Multi-Grid Methods, volume 11 of Notes on Numerical Fluid Mechanics, pages 33–44. Friedrich Vieweg und Sohn, Braunschweig, Germany, 1984.

Hemker:1985:MSSa

[HS85a] P. W. Hemker and S. P. Spekreijse. Multigrid solution of the steady Euler equations. In Advances in multigrid methods (Oberwolfach, 1984), volume 11 of Notes Numer. Fluid Mech., pages 33–44. Friedrich Vieweg und Sohn, Braunschweig, Germany, 1985.

Hemker:1985:MSSb

[HS85b] P. W. Hemker and S. P. Spekreijse. Multigrid solution of the steady Euler equations. CWI Reports NM, R8505. Stichting Mathematisch Centrum, Centrum voor Wiskunde en Informatica, Amsterdam, The Netherlands, 1985. 13 pp.

Hemker:1986:MGO

[HS86] P. W. Hemker and S. P. Spekreijse. Multiple grid and Osher’s scheme for the efficient solution of the steady Euler equations. Applied Numerical Mathematics: Transactions of IMACS, 2(6):475–493, December 1986. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Hemker:1992:DAS

[HS92] P. W. Hemker and G. I.

Shishkin. Discrete approximation of singularly perturbed parabolic PDEs with a discontinuous initial condition. In J. J. H. Miller, editor, BAIL VI Proceedings, pages 3–4. Front Range Press, Copper Mountain, Colorado, USA, 1992.

Hemker:1993:APP

[HS93] P. W. Hemker and G. I.

Shishkin. Approximation of parabolic PDEs with a discontinuous initial condition. East-West Journal of Numerical Mathematics, 1(4):287–302, 1993. CODEN EJNMEA. ISSN 0928-0200.

Hemker:1994:DAS

[HS94a] P. W. Hemker and G. I.

Shishkin. Discrete approximation of singularly perturbed parabolic PDEs with a discontinuous initial condition. Computational Fluid Dynamics Journal, 2:375–392, 1994.

Hemker:1994:CSP

[HS94b] P. W. Hemker and G. I.

Shishkin. On a class of singularly perturbed boundary value problems for which an adaptive mesh technique is necessary. In D. Bainov and V. Covachev, editors, Proceedings of the Second International Colloquium on Numerical Analysis (Plovdiv, August 13–17, 1993), pages 83–92. VSP, International Science Publishers, Utrecht, The Netherlands, 1994. ISBN 90-6764-152-9. Proceedings of the Second International Colloquium on Numerical Analysis.

Hemker:1995:UDC

[HS95] P. W. Hemker and G. I. Shishkin. The use of defect correction for the solution of parabolic singular perturbation problems. In Abstracts of the International Conference on Advanced Mathematics, V. Computations and Applications, pages 135–136. ????, ????, 1995. Novosibirsk, June 20–24.

Hemker:2001:RFF

[HS01] P. W. Hemker and F. Sprengel. On the representation of functions and finite difference operators on adaptive dyadic grids. Computational Methods in Applied Mathematics, 1(3): 222–241, 2001. ISSN 1609-4840 (print), 1609-9389 (electronic).

Hemker:1980:CSA

[HSd80] P. W. Hemker, H. Schippers, and P. M. de Zeeuw. Comparing some aspects of two codes for two-point boundary-value problems. Technical Report NW 98, Mathematical Centre, Amsterdam, The Netherlands, 1980.

Hemker:1997:USH

[HSS97a] P. W. Hemker, G. I. Shishkin, and L. P. Shishkina. ϵ-uniform schemes with high-order timeaccuracy for parabolic singular perturbation problems. IMA Journal of Numerical Analysis, submitted, 1997. CO-DEN IJNADH. ISSN 0272-4979 (print), 1464-3642 (electronic).

Hemker:1997:DMA

[HSS97b] P. W. Hemker, G. I. Shishkin, and L. P. Shishkina. On discrete methods arising in the numerical solution of singularly perturbed convection-diffusion problems, 1997.

Hemker:1997:UDC

[HSS97c] P. W. Hemker, G. I. Shishkin, and L. P. Shishkina. The use of defect correction for the solution of parabolic singular perturbation problems. Zeitschrift für Angewandte Mathematik und Mechanik, 76(1):59–74, 1997. CODEN ZAMMAX. ISSN 0044-2267 (print), 1521-4001 (electronic).

Hemker:1998:DCM

[HSS98] Pieter W. Hemker, Grigorii I.

Shishkin, and Lidia P. Shishkina. A defect correction method for parabolic singular perturbation problems on a rectangle. In Second International Conference “Finite-Difference Methods: Theory and Application”, Vol. 2 (Minsk, 1998), pages 30–34. Natl. Acad. Sci. Belarus Inst. Math., Minsk, 1998.

Hemker:1999:UDC

[HSS99a] P. W. Hemker, G. I. Shishkin, and L. P. Shishkina. An ϵ-uniform defect-correction method for a parabolic convectiondiffusion problem. In Recent advances in numerical methods and applications, II (Sofia, 1998), pages 521–529. World Sci. Publ., River Edge, NJ, 1999.

Hemker:1999:NSN

[HSS99b] P. W. Hemker, G. I. Shishkin, and L. P. Shishkina. The numerical solution of a Neumann problem for parabolic singularly perturbed equations with high-order time accuracy. In Recent advances in numerical methods and applications, II (Sofia, 1998), pages 27–39. World Sci. Publ., River Edge, NJ, 1999.

Hemker:2000:USH

[HSS00] P. W. Hemker, G. I. Shishkin, and L. P. Shishkina. ϵ-uniform schemes with high-order timeaccuracy for parabolic singular perturbation problems. IMA Journal of Numerical Analysis, 20(1):99–121, January 2000. CODEN IJ-NADH. ISSN 0272-4979 (print), 1464-3642 (electronic).

URL http://www3.oup.co.
uk/imanum/hdb/Volume_20/
Issue_01/200099.sgm.abs.
html;   http://www3.oup.
co.uk/imanum/hdb/Volume_
20/Issue_01/pdf/200099.pdf.

Hemker:2001:HOT

[HSS01a] P. W. Hemker, G. I. Shishkin, and L. P. Shishkina. Highorder time-accurate parallel schemes for parabolic singularly perturbed problems with convection. Computing: Archiv für informatik und numerik, 66(2):139–161, 2001. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL

http://link.springer-ny.
com/link/service/journals/
00607/bibs/1066002/10660139.
htm;   http://link.springer-
ny.com/link/service/journals/
00607/papers/1066002/10660139.

pdf. Archives for scientific computing. Numerical methods for transport-dominated and related problems (Magdeburg, 1999).

Hemker:2001:APC

[HSS01b] Pieter W. Hemker, Grigorii I.

Shishkin, and Lidia P. Shishkina. Acceleration by parallel computations of solving high-order time-accurate difference schemes for singularly perturbed convectiondiffusion problems. In Numerical analysis and its applications (Rousse, 2000), volume 1988 of Lecture Notes in Comput. Sci., pages 393–401. Springer-Verlag, Berlin, Germany /Heidelberg, Germany / London, UK / etc., 2001.

Hemker:2002:DCH

[HSS02a] P. W. Hemker, G. I. Shishkin, and L. P. Shishkina. Defect correction high-order accurate schemes for parabolic singularly perturbed convectiondiffusion problems. In International Conference on Computational Mathematics. Part I, II, pages 454–459. ICM&MG Publishing, Novosibirsk, Russia, 2002.

Hemker:2002:HOTb

[HSS02b] P. W. Hemker, G. I. Shishkin, and L. P. Shishkina. Highorder time-accuracy schemes for parabolic singular perturbation problems with convection. Russian journal of numerical analysis and mathematical modelling, 17(1):1–24, 2002. CODEN RJNMEH. ISSN 0927-6467 (print), 1569-3988 (electronic).

Hemker:2002:HOTa

[HSS02c] Pieter W. Hemker, Grigorii I. Shishkin, and Lidia P. Shishkina. High-order time-accurate schemes for singularly perturbed parabolic convectiondiffusion problems with Robin boundary conditions. Computational Methods in Applied Mathematics, 2(1):3–25, 2002. ISSN 1609-4840 (print), 1609-9389 (electronic).

Hemker:2003:HOT

[HSS03a] P. W. Hemker, G. I. Shishkin, and L. P. Shishkina. Highorder time-accurate schemes for parabolic singular perturbation convection-diffusion problems with Robin boundary conditions. Mat. Model., 15(8):99– 112, 2003. ISSN 0234-0879.

Hemker:2003:NDC

[HSS03b] Pieter W. Hemker, Grigorii I. Shishkin, and Lidia P. Shishkina. Novel directcorrection high-order, in space and time, accurate schemes for parabolic singularly perturbed convection-diffusion problems. Computational Methods in Applied Mathematics, 3(3):387– 404 (electronic), 2003. ISSN 1609-4840 (print), 1609-9389 (electronic). Dedicated to John J. H. Miller on the occasion of his 65th birthday.

Hemker:2002:PUN

[HST02] P. W. Hemker, G. I. Shishkin, and I. V. Tselishcheva. Parameteruniform numerical methods for a class of quasilinear singularly perturbed convectiondiffusion equations in a composed domain. In International Conference on Computational Mathematics. Part I, II, pages 460–466. ICM&MG Publishing, Novosibirsk, Russia, 2002.

Hemker:1990:BDS

[HvdME90] P. W. Hemker, H. T. M.

van der Maarel, and C. T. H. Everaars. BASIS: A data structure for adaptive multigrid computations. Technical Report NM-R9014, Centrum voor Wiskunde en Informatica, P. O. Box 4079, 1009 AB Amsterdam, The Netherlands, 1990.

Hemker:2004:FTLa

[HvR04] P. W. Hemker and M. H. van Raalte. Fourier two-level analysis for higher dimensional discontinuous Galerkin discretisation. Computing and Visualization in Science, 7(3–4):159– 172, 2004. ISSN 1432-9360 (print), 1433-0369 (electronic).

Hemker:1979:PRN

[HW79] P. W. Hemker and D. T. Winter. A preliminary report on numerical operators in ALGOL 68. Technical Report NW 66, Mathematical Centre, Amsterdam, The Netherlands, 1979.

Hemker:1994:CM

[HW94a] P. W. Hemker and P. Wesseling, editors. Contributions to Multigrid, volume 103 of CWI Tract Series. Centrum voor Wiskunde en Informatica, P. O. Box 4079, 1009 AB Amsterdam, The Netherlands, 1994.

Hemker:1994:MMI

[HW94b] P. W. Hemker and P. Wesseling, editors. Multigrid methods IV, volume 116 of International Series of Numerical Mathematics. Birkhäuser Verlag, Basel, 1994. Proceedings of the Fourth European Multigrid Conference, held in Amsterdam, July 6–9, 1993.

Hemker:1984:PVCa

[HWD84] P. W. Hemker, P. Wesseling, and P. M. De Zeeuw. A portable vector-code for autonomous multigrid modules. In B. Enquist and T. Smedsass, editors, Proc. IFIP-Conference on PDE Software: Modules, Interfaces and Systems, pages 29–40. North-Holland, Amsterdam, The Netherlands, 1984.

Koren:1988:LGS

[KdH88] B. Koren, P. M. de Zeeuw, and P. W. Hemker. Line Gauss– Seidel relaxation and multigrid for steady, two-dimensional flow computations. Technical Report NM-N8803, Centrum voor Wiskunde en Informatica, P. O. Box 4079, 1009 AB Amsterdam, The Netherlands, 1988.

Koren:1991:DDDa

[KH91a] B. Koren and P. W. Hemker.

Damped, direction dependent multigrid for hypersonic flow computations. Applied Numerical Mathematics: Transactions of IMACS, 7(4):309– 328, 1991. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Koren:1991:EMC

[KH91b] B. Koren and P. W. Hemker.

Efficient multigrid computation of steady hypersonic flows. In Computational methods in hypersonic aerodynamics, volume 9 of Fluid Mech. Appl., pages 203–231. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1991.

Koren:1991:MDUa

[KH91c] B. Koren and P. W. Hemker. Multi-dimensional upwind schemes, multigrid and defect correction for accurate and efficient Euler flow computations. In W. Berry and B. Battrick, editors, Aerodynamics for Space Vehicles, pages 129–134. European Space Agency, ESTEC, Noordwijk, The Netherlands, 1991.

Koren:1992:EMC

[KH92a] B. Koren and P. W. Hemker. Efficient multigrid computation of steady hypersonic flows. In T. K. S. Murthy, editor, Computational Methods in Hypersonic Aerodynamics, pages 203–231. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1992.

Koren:1992:MUM

[KH92b] B. Koren and P. W. Hemker. Multi-D upwinding and multigridding for steady Euler flow computations. In J. B. Vos, A. Rizzi, and I. L. Ryhming, editors, Proceedings of the Ninth GAMM-Conference on Numerical Methods in Fluid Mechanics (Lausanne, 1991), volume 35 of Notes on Numerical Methods in Fluid Mechanics, pages 89–98. Friedrich Vieweg und Sohn, Braunschweig, Germany, 1992.

Koren:1997:GTS

[KH97] B. Koren and P. W. Hemker.

Godunov-type schemes, sparse grids and 3D CFD. Journal of computational physics, 137(??), ???? 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). To appear.

Koren:1997:SCT

[KHd97] B. Koren, P. W. Hemker, and P. M. de Zeeuw. Semicoarsening in three directions for Euler-flow computations in three dimensions. In Euler and Navier–Stokes Solvers Using Multi-Dimensional Upwind Schemes and Multigrid Acceleration, volume 57 of Notes on Numerical Fluid Mechanics, pages 547–567. Friedrich Vieweg und Sohn, Braunschweig, Germany, 1997.

Koren:1997:MSC

[KHE97] B. Koren, P. W. Hemker, and C. T. H. Everaars. Multiple semi-coarsened multigrid for 3D CFD. AIAA paper, 97(2029), 1997. CODEN AAPRAQ. ISSN 0146-3705.

Khemker:2004:HOA

[KSS04] P. V. Khemker, G. I. Shishkin, and L. P. Shishkina. Highorder accuracy decomposition of Richardson’s method for a singularly perturbed elliptic reaction-diffusion equation. Zh. Vychisl. Mat. Mat. Fiz., 44(2):

329–337, 2004. ISSN 0044-4669.

Koren:2003:FSE

[KvBH^{+}03] B. Koren, E. H. van Brummelen, P. W. Hemker, B. van Leer, and M. R. Lewis. Fix for solution errors near interfaces in two-fluid flow computations. In Computational fluid dynamics, 2002 (Sydney), pages 523–528. Springer-Verlag, Berlin, Germany /Heidelberg, Germany / London, UK / etc., 2003.

Molenaar:1990:MASa

[MH90] J. Molenaar and P. W. Hemker. A multigrid approach for the solution of the 2D semiconductor equations. Impact Comput. Sci. Eng., 2:219–243, 1990. CODEN ICOEEK. ISSN 0899-8248 (print), 1557-7678 (electronic).

Noordmans:1997:CRS

[NH97] J. Noordmans and P. W. Hemker. Convergence results for 3D sparse grid approaches. In PRISM’97, Proceedings of the Conference on Preconditioned Iterative Solution Methods for Large Scale Problems in Scientific Computations, pages 11–22. Nijmegen University, Nijmegen, The Netherlands, 1997.

Noordmans:1998:CRS

[NH98] J. Noordmans and P. W. Hemker. Convergence results for 3D sparse grid approaches. Numerical Linear Algebra with Applications, 5 (5):363–376 (1999), September/October 1998. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). Special Issue: PRISM 97 (Nijmegen).

Noordmans:2000:AAS

[NH00] J. Noordmans and P. W.

Hemker. Application of an adaptive sparse-grid technique to a model singular perturbation problem. Computing: Archiv für informatik und numerik, 65(4):357–378, 2000. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Polak:1981:MPC

[PWBH81] S. J. Polak, A. Wachters, Th.

Beelen, and P. W. Hemker. A mesh-parameter continuation method. In M. H. Schultz, editor, Elliptic Problem Solvers, pages 383–390. Academic Press, New York, NY, USA, 1981.

Rispens:2011:LFA

[RTH^{+}11] Theo Rispens, Henk Te Velthuis, Piet Hemker, Han Speijer, Wim Hermens, and Lucien Aarden. Label-free assessment of high-affinity antibody-antigen binding constants. Comparison of bioassay, SPR, and PEIA-ellipsometry. Journal of Immunological Methods, 365(1–2):50–57, February 2011. ISSN 0022-1759.

Shishkin:1997:NMF

[SSH97] G. I. Shishkin, L. P. Shishkina, and P. W. Hemker. A numerical method with floating meshes for singularly perturbed problems with a concentrated disturbance in the initial data. CWI Quarterly, 10(3-4):317–336, 1997. ISSN 0922-5366. International Workshop on the Numerical Solution of Thin-layer Phenomena (Amsterdam, 1997).

Shishkin:2004:CSP

[SSH04] Grigory I. Shishkin, Lidia P. Shishkina, and Pieter W. Hemker. A class of singularly perturbed convection-diffusion problems with a moving interior layer. An a posteriori adaptive mesh technique. Computational Methods in Applied Mathematics, 4(1):105– 127 (electronic), 2004. ISSN 1609-4840 (print), 1609-9389 (electronic).

Thoolen:1991:AMC

[TH91] P. M. C. Thoolen and P. W. Hemker. Approximation methods for n-component solute transport and ion-exchange. Technical Report Report 91-09, Dept. of Mathematics, Univ. of Amsterdam, 1991.

Thoolen:1994:AMC

[TH94] P. M. C. Thoolen and P. W. Hemker. Approximate methods for n-component solute transport and ion-exchange.

Journal of Computational and Applied Mathematics, 53(2): 275–290, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Maarel:1990:EAE

[vdMHE90] H. T. M. van der Maarel, P. W.

Hemker, and C. T. H. Everaars. EULER: An adaptive Euler code. Technical Report NM-R9015, Centrum voor Wiskunde en Informatica, P. O. Box 4079, 1009 AB Amsterdam, The Netherlands, 1990.

vanderMaarel:1993:ASA

[vdMHKM93] H. T. M. van der Maarel, P. W. Hemker, B. Koren, and J. A. Michelsen. Application of a solution-adaptive multigrid method to the Euler equations. CWI Quarterly, 6(1):49– 75, 1993. ISSN 0922-5366.

Domselaar:1975:NPE

[vH75] B. van Domselaar and P. W.

Hemker. Nonlinear parameter estimation in initial value problems. Technical Report NW 18, MC, 1975.

VanderMaarel:1995:SAF

[VH95] H. T. M. Van der Maarel and P. W. Hemker. Structured adaptive finite-volume multigrid for compressible flows. In W. Hackbusch and G. Wittum, editors, Fast Solvers for Flow Problems (Kiel, 1994), volume 49 of Notes on Numerical Fluid Mechanics, pages 14– 37. Friedrich Vieweg und Sohn, Braunschweig, Germany, 1995.

Vasileva:2006:AMS

[VKH06] Daniela Vasileva, Anton Kuut, and Pieter W. Hemker. An adaptive multigrid strategy for convection-diffusion problems. In Large-scale scientific computing, volume 3743 of Lecture Notes in Comput. Sci., pages 138–145. Springer-Verlag, Berlin, Germany /Heidelberg, Germany / London, UK / etc., 2006.

vanRaalte:2005:TLM

[vRH05] M. H. van Raalte and P. W. Hemker. Two-level multigrid analysis for the convectiondiffusion equation discretized by a discontinuous Galerkin method. Numerical Linear Algebra with Applications, 12(5– 6):563–584, June/August 2005. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Wagenvoord:2006:LSC

[WHH06] R. Wagenvoord, P. W. Hemker, and H. C. Hemker. The limits of simulation of the clotting system. Journal of Thrombosis and Haemostasis: JTH, 4(6): 1331–1338, June 2006. ISSN 1538-7836 (print), 1538-7933 (electronic).

Wesseling:2008:ESI

[WHO08] Pieter Wesseling, Pieter Hemker, and Cornelis Oosterlee. Editorial [Special issue devoted to the 8th European Multigrid Conference (EMG)]. Computing and Visualization in Science, 11(1):1, 2008. ISSN 1432-9360 (print), 1433-0369 (electronic). Held in Scheveningen, September 27–30, 2005.